ICM 1998, Berlin, Aug. 18–27 Abstracts of Plenary and Invited Lectures

Section:

6. Topology

1991 MS Classification: 57R55 Stern, Ronald J., University of California at Irvine, USA Knot Surgery on 4-Manifolds (joint talk with R. Fintushel)

ICM 1998 Berlin · Plenary and Invited Lectures